# A Bimodal Metallicity Distribution Function in the Ultra-Faint Dwarf Galaxy Reticulum II

### v1

@@ 6 CONCLUSION
The star formation histories of ultra-faint satellites of the Milky Way have revealed that most of their stars formed in the early Universe before , making UFDs extremely ancient systems (Brown et al., 2014; Jenkins et al., 2021; Simon et al., 2023). However, studies suggest that there is star formation after reionization highlighting the importance of the environment in fully quenching these systems (Sacchi et al., 2021; Savino et al., 2023; McQuinn et al., 2023). To study the star formation histories of UFDs, we need well-populated MDFs. The MDFs for large dwarf galaxies around the Milky Way have been well studied (Kirby et al., 2011), however, UFDs are more difficult to get well-populated spectroscopic MDFs and are not as extensively studied.
In this work, we increased the number of spectroscopic metallicities in the Reticulum II UFD by 6.5 times. This is the most populated MDF of any UFD. We detect a bimodal metallicity distribution with a low metallicity peak at with about of the stars and a high metallicity peak at with about of the stars. This is the first bimodal MDF in any UFD. The bimodal MDF is consistent with the current two-burst star formation history in Ret II, which forms of its stars in one burst during the early universe and of its stars a few Gyrs later, after reionization (Simon et al., 2023). The age-metallicity relation from the two-burst SFH in the literature and our bimodal MDF relates the metallicity gap of dex to the 3 Gyr age gap, suggesting that there is star formation after reionization. However, we need to get stellar ages using our metallicities to be able to distinguish this scenario from a multiple progenitor halo accretion scenario. Further, with more well-populated UFD MDFs, we will be able to investigate the relative importance between reionization and internal quenching on the smallest scale galaxies.

### v2

@@ 6 CONCLUSION
The star formation histories of ultra-faint satellites of the Milky Way have revealed that most of their stars formed in the early Universe before , making UFDs extremely ancient systems (Brown2014; Jenkins2021; Simon+2023). However, studies suggest that there is star formation after reionization highlighting the importance of the environment in fully quenching these systems (Sacchi+2021; Savino2023; McQuinn2023_PegasusW). To study the star formation histories of UFDs, we need well-populated MDFs. The MDFs for large dwarf galaxies around the Milky Way have been well studied (Kirby+2011), however, UFDs are more difficult to get well-populated spectroscopic MDFs and are not as extensively studied.
In this work, we increased the number of spectroscopic metallicities in the Reticulum II UFD by 6.5 times. This is the most populated MDF of any UFD. We detect a bimodal metallicity distribution with a low metallicity peak at with about of the stars and a high metallicity peak at with about of the stars. This is the first bimodal MDF in any UFD. The bimodal MDF is consistent with the current two-burst star formation history in Ret II, which forms of its stars in one burst during the early universe and of its stars a few Gyrs later, after reionization (Simon+2023). The age-metallicity relation from the two-burst SFH in the literature and our bimodal MDF relates the metallicity gap of dex to the 3 Gyr age gap, suggesting that there is star formation after reionization. However, we need to get stellar ages using our metallicities to be able to distinguish this scenario from a multiple progenitor halo accretion scenario. Further, with more well-populated UFD MDFs, we will be able to investigate the relative importance between reionization and internal quenching on the smallest scale galaxies.